# Multidimensional empirical wavelet transform

### v1

@@ 4.2.2 Non-compactly supported $\boldsymbol{\widehat{\psi}}$
In this section, we assume that the support $\overline{\Lambda}$ of $\widehat{\psi}$ is not compact. The following theorems state sufficient conditions for which the system
$\{T_{b_{n}k}\psi_{n}\}_{(n,k)\in\Upsilon\times{\mathbb{Z}}^{N}}\quad\textrm{ and }\quad\{T_{b_{n}k}\phi_{n}\}_{(n,k)\in\Upsilon^{+}\times{\mathbb{Z}}^{N}}$
@@
$=|b_{n}|^{-1}\sum_{k\in{\mathbb{Z}}^{N}}\left|\int_{b_{n}^{-1}\mathbb{T}}\left(\sum_{l\in{\mathbb{Z}}^{N}}\widehat{f}(\xi-b_{n}^{-1}l)\overline{\widehat{\psi}_{n}(\xi-b_{n}^{-1}l)}\right)e^{2\pi\imath\frac{1}{b_{n}^{-1}}\left(k\cdot\xi\right)}{\mathrm{d}}\xi\right|^{2},$
@@
$A=\inf_{\xi\in{\mathbb{R}}^{N}}\left(\sum_{n\in\Upsilon^{+}}\frac{1}{|b_{n}|}\left|\hat{\phi}_{n}(\xi)\right|^{2}-\sum_{n\in\Upsilon^{+}}\sum_{k\neq 0}\frac{1}{|b_{n}|}\left|\hat{\phi}_{n}(\xi)\widehat{\phi}_{n}(\xi-b_{n}^{-1}k)\right|\right)>0$
@@
$B=\sup_{\xi\in{\mathbb{R}}^{N}}\sum_{n\in\Upsilon^{+}}\sum_{k\in\mathbb{Z}^{N}}\frac{1}{|b_{n}|}\left|\hat{\phi}_{n}(\xi)\widehat{\phi}_{n}(\xi-b_{n}^{-1}k)\right|<\infty,$
then the system $\{T_{b_{n}k}\phi_{n}\}_{n\in\Upsilon^{+},k\in\mathbb{Z}^{N}}$ is a frame for $\mathrm{L}^{2}({\mathbb{R}}^{N})$ with frame bounds $A$ and $B$.
@@
This results from the argumentation of the proof of Theorem 5 by replacing $\Upsilon$ by $\Upsilon^{+}$ and $\psi_{n}$ by $\phi_{n}$. ∎
@@ 5 Empirical wavelet systems from affine deformations
In this section, we consider a partition $\{\Omega_{n}\}_{n\in\Upsilon}$ for which we can find diffeomorphisms $\varphi_{n}$ that are affine functions, i.e., of the form $\varphi_{n}(\xi)=\beta_{n}(\xi-\eta_{n})$ with $\beta_{n}$ a linear function. If the construction of the empirical wavelet systems introduced in Section 3 is natural in the Fourier domain, it is also possible to build them directly in the spatial domain for affine mappings.
@@
Let $\beta$ be a linear function (non identically zero). The function $g$ corresponding to the inverse Fourier transform of the deformation of the Fourier transform $\widehat{f}$ of a function $f$ by $\beta$, i.e., $\widehat{g}=\widehat{f}\circ\beta$, is given by, for every $x\in{\mathbb{R}}^{N}$,
$g(x)=\left(f\star{\mathcal{F}}^{-1}\left(\left|det\;J_{\beta^{-1}}\right|\right)\right)(\beta^{-\intercal}(x)),$ (12)
@@
$g(x)=\int_{{\mathbb{R}}^{N}}\widehat{f}(\beta(\xi))e^{2\pi\imath(\xi\cdot x)}\mathrm{d}\xi$ $=\int_{\phi({\mathbb{R}}^{N})}\widehat{f}(u)e^{2\pi\imath(\beta^{-1}(u)\cdot x)}\left|\mathrm{det}\;J_{\beta^{-1}}(u)\right|\mathrm{d}u$
$=\int_{\phi({\mathbb{R}}^{N})}\widehat{f}(u)e^{2\pi\imath(u\cdot\beta^{-\intercal}(x))}\left|\mathrm{det}\;J_{\beta^{-1}}(u)\right|\mathrm{d}u$
$={\mathcal{F}}^{-1}\left(\widehat{f}\bullet\left|\mathrm{det}\;J_{\beta^{-1}}\right|\right)(\beta^{-\intercal}(x))$
$=\left(f\star{\mathcal{F}}^{-1}\left(\left|\mathrm{det}\;J_{\beta^{-1}}\right|\right)\right)(\beta^{-\intercal}(x)).$
@@
Let $\{\Omega_{n}\}_{n\in\Upsilon}$ be a partition of the Fourier domain and $\psi$ a mother wavelet. Let assume there exists a set of linear functions $\{\beta_{n}\}_{n\in\Upsilon}$ such that $\Lambda=\beta_{n}(\Omega_{n}-\eta_{n})$ if $\Omega_{n}$ is bounded and $\Lambda\subsetneq\beta_{n}(\Omega_{n}-\eta_{n})$ otherwise. The set of empirical wavelets, $\{\psi_{n}\}_{n\in\Upsilon}$, defined in the Fourier domain by, for every $\xi\in{\mathbb{R}}^{N}$,
$\widehat{\psi}_{n}(\xi)=\sqrt{\left|\mathrm{det}J_{\beta_{n}}(\xi-\eta_{n})\right|}\,\widehat{\psi}(\beta_{n}(\xi-\eta_{n})),$
@@
$\psi_{n}(x)=\left(\psi\star{\mathcal{F}}^{-1}\left(\sqrt{|\mathrm{det}\;J_{\beta_{n}^{-1}}|}\right)\right)(\beta_{n}^{-\intercal}(x))\,e^{2\pi\imath\,\eta_{n}x}.$
@@
First, since the inverse of $J_{\beta_{n}^{-1}}(\beta_{n}(\xi))$ is given by $J_{\beta_{n}}(\xi)$, we can rewrite
$\sqrt{|\mathrm{det}\;J_{\beta_{n}}(\xi-\eta_{n})|}=\frac{1}{\sqrt{\left|\mathrm{det}\;J_{\beta^{-1}_{n}}(\beta_{n}(\xi-\eta_{n}))\right|}}.$
@@
$\psi_{n}(x)$ $={\mathcal{F}}^{-1}\left(\left(\frac{1}{\sqrt{|\mathrm{det}J_{\beta_{n}^{-1}}|}}\bullet\widehat{\psi}\right)\circ\beta_{n}\right)(x)\,e^{2\pi\imath\eta_{n}x}.$
By applying Lemma 1 with $\widehat{f}=\frac{1}{\sqrt{|\mathrm{det}J_{\beta_{n}^{-1}}|}}\bullet\widehat{\psi}$ and $\beta=\beta_{n}$, we obtain
$\psi_{n}(x)$ $=\left({\mathcal{F}}^{-1}\left(\frac{1}{\sqrt{|\mathrm{det}J_{\beta_{n}^{-1}}|}}\bullet\widehat{\psi}\right)\star{\mathcal{F}}^{-1}\left(|\mathrm{det}\;J_{\beta_{n}^{-1}}|\right)\right)(\beta_{n}^{-\intercal}(x))\,e^{2\pi\imath\,\eta_{n}x}$
$=\left({\mathcal{F}}^{-1}\left(\frac{1}{\sqrt{|\mathrm{det}J_{\beta_{n}^{-1}}|}}\bullet\widehat{\psi}\bullet|\mathrm{det}\;J_{\beta_{n}^{-1}}|\right)\right)(\beta_{n}^{-\intercal}(x))\,e^{2\pi\imath\,\eta_{n}x}$
$=\left({\mathcal{F}}^{-1}\left(\widehat{\psi}\bullet\sqrt{|\mathrm{det}\;J_{\beta_{n}^{-1}}|}\right)\right)(\beta_{n}^{-\intercal}(x))\,e^{2\pi\imath\,\eta_{n}x}$
$=\left(\psi\star{\mathcal{F}}^{-1}\left(\sqrt{|\mathrm{det}\;J_{\beta_{n}^{-1}}|}\right)\right)(\beta_{n}^{-\intercal}(x))\,e^{2\pi\imath\,\eta_{n}x}.$
@@
As in Example 3.1, we consider the 1D case ($N=1$) with $\widehat{\psi}$ of support $\Lambda$ the open interval of size $1$ centered on 00 and $\{\Omega_{n}\}_{n\in\Upsilon}$ a partition of open bounded intervals with center $\omega_{n}$. The diffeomorphism $\varphi_{n}:\xi\mapsto\frac{1}{|\Omega_{n}|}(\xi-\omega_{n})$, verifying $\Lambda=\varphi_{n}(\Omega_{n})$, can be rewritten $\varphi_{n}(\xi)=\beta_{n}(\xi-\eta_{n})$ with $\beta_{n}:\xi\mapsto\xi/|\Omega_{n}|$ and $\eta_{n}=\omega_{n}$. The associated empirical wavelet system is therefore given in the spatial domain by, for every $x\in{\mathbb{R}}$,
$\psi_{n}(x)$ $=\left(\psi\star{\mathcal{F}}^{-1}\left(\sqrt{|\Omega_{n}|}\right)\right)(|\Omega_{n}|x)\,e^{2\pi\imath\,\omega_{n}x}$ $=\sqrt{|\Omega_{n}|}\,\psi(|\Omega_{n}|x)\,e^{2\pi\imath\,\omega_{n}x}.$
@@
The construction of a symmetric empirical wavelet system $\{\phi_{n}\}_{n\in\Upsilon}$, induced by affine diffeomorphisms, in the spatial domain stems from Proposition 2, using Equation 3.
@@
We consider again the 1D domain. From the system $\{\psi_{n}\}_{n\in\Upsilon}$ defined as in Example 5.1, we can build a symmetric system $\{\phi_{n}\}_{n\in\Upsilon}$ as follows, for every $x\in{\mathbb{R}}$,
$\phi_{0}(x)=\sqrt{|\Omega_{0}|}\,\psi(|\Omega_{0}|x),$
@@
$\phi_{n}(x)$ $=\sqrt{\frac{|\Omega_{n}|}{2}}\psi(|\Omega_{n}|x)e^{2\pi\imath\,\omega_{n}x}+\sqrt{\frac{|\Omega_{n}|}{2}}\psi(|\Omega_{n}|x)e^{-2\pi\imath\,\omega_{n}x}$
@@ 6 Construction of empirical wavelet systems
In this section, 2D continuous empirical wavelet systems are constructed from Gabor and Shannon mother wavelets and we examine their guarantees of reconstruction. Their practical behaviors are analyzed through numerical experiments conducted on two different (real-valued) images: a toy image of size $256\times 256$ and the classic Barbara image of size $512\times 512$. These images are shown in Figure 3 (left).
@@
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@
To numerically construct empirical wavelet systems $\{\phi_{n}\}_{n\in\Upsilon^{+}}$ as in Equation 2, we need to compute an estimate $\breve{\varphi}_{n}$ of the diffeomorphism $\varphi_{n}$. This estimation is performed using the Demons algorithm [27]. For each region $\Omega_{n}$, the pair of smoothing parameter and number of multiresolution image pyramid levels is selected by a grid search minimizing the quadratic risk $\|\Lambda-\breve{\varphi}_{n}(\Omega_{n})\|_{2}^{2}$ on the values in $(0.3,0.35,\ldots,0.7)\times(n_{P}-2,n_{P}-1,n_{P})$, where $n_{P}$ is the highest integer such that $2^{n_{P}}$ is smaller than each dimension of the image. The number of iterations at the $n_{P}$ pyramid levels are set to $(2^{4},\ldots,2^{n_{P}+1})$ from the highest to the lowest pyramid level.
For all the numerical experiments, the symmetric empirical wavelet transform is computed as in Definition 6 with $b_{n}=1$ for every $n\in\Upsilon^{+}$, so that the results for continuous wavelet frames in Section 4 apply. To each wavelet coefficient ${\mathcal{E}}_{\phi}^{f}(\cdot,n)$ of an image $f$ of size $M\times M$, we associate the wavelet spectrum defined as $|{\mathcal{E}}_{\phi}^{f}(\cdot,n)|^{2}$ and the Fourier spectrum energy $\mathrm{E}_{n}$ of the underlying region $\Omega_{n}\cup\Omega_{-n}$ that reads
@@
The Fourier transform of the 1D Gabor mother wavelet is given by (see [7]), for every $u\in{\mathbb{R}}$,
@@
Let assume that the diffeomorphisms $\varphi_{n}$ satisfy, for $a.e.\,\xi\in{\mathbb{R}}^{2}$, |$|\{m\in\Upsilon\mid\varphi_{m}(\xi)=\varphi_{n}(\xi)\}|\leq K_{\xi}\in\mathbb{N}$ for every $n\in\Upsilon$ and $\{|\mathrm{det}\;J_{\varphi_{n}}(\xi)|\}_{n\in\Upsilon}$ is a bounded sequence. Then, the continuous reconstruction is guaranteed for the systems $\{\psi_{n}\}_{n\in\Upsilon}$ and $\{\phi_{n}\}_{n\in\Upsilon}$ induced by $\widehat{\psi}^{\mathrm{G}}\circ\varphi_{n}$, and is given by Equations 8 and 9.
@@
$\sum_{n\in\Upsilon}\left|\widehat{\psi}_{n}(\xi)\right|^{2}=\sum_{n\in\Upsilon}|\mathrm{det}\;J_{\varphi_{n}}(\xi)|\left|\widehat{\psi}(\varphi_{n}(\xi))\right|^{2}=\sum_{n\in\Upsilon}|\mathrm{det}\;J_{\varphi_{n}}(\xi)|e^{-2\pi(\frac{5}{2})^{2}\|\varphi_{n}(\xi)\|^{2}}.$
First, since $\varphi_{n}$ is a diffeomorphism, we have, for every $\xi\in{\mathbb{R}}^{2}$, $|\mathrm{det}\;J_{\varphi_{n}}(\xi)|>0$ and therefore
@@
Now, let us denote $\Gamma_{\xi}=\left\{m\in\Upsilon\mid\varphi_{m}(\xi)=\varphi_{n}(\xi)\Rightarrow n\geq m\right\}$. For $a.e.\,\xi\in{\mathbb{R}}^{2}$, the $\varphi_{n}(\xi)$ for $n\in\Gamma_{\xi}$ are all different and the condition |$|\{m\in\Upsilon\mid\varphi_{m}(\xi)=\varphi_{n}(\xi)\}|\leq K_{\xi}\in\mathbb{N}$, for every $n\in\Upsilon$, means that there are at most $K_{\xi}$ integers $n$ for which the $\varphi_{n}(\xi)$ have the same value. It follows that, for $a.e.\,\xi\in{\mathbb{R}}^{2}$,
$\sum_{n\in\Upsilon}\left|\widehat{\psi}_{n}(\xi)\right|^{2}$ $\leq\max_{n\in\Upsilon}|\mathrm{det}\;J_{\varphi_{n}}(\xi)|\sum_{n\in\Upsilon}e^{-2\pi(\frac{5}{2})^{2}\|\varphi_{n}(\xi)\|^{2}}$
$\leq K_{\xi}\max_{n\in\Upsilon}|\mathrm{det}\;J_{\varphi_{n}}(\xi)|\sum_{n\in\Gamma_{\xi}}e^{-2\pi(\frac{5}{2})^{2}\|\varphi_{n}(\xi)\|^{2}}$
$\leq K_{\xi}\max_{n\in\Upsilon}|\mathrm{det}\;J_{\varphi_{n}}(\xi)|\int_{{\mathbb{R}}^{2}}e^{-2\pi(\frac{5}{2})^{2}\|u\|^{2}}\mathrm{d}u<\infty.$
Then, Theorem 1 applies and gives a dual frame of $\{\psi_{n}\}_{n\in\Upsilon}$. This guarantees the reconstruction using Equation 8.
@@
$\sum_{n\in\Upsilon\setminus\{0\}}\left|\widehat{\psi}_{n}(\xi)\right|\left|\widehat{\psi}_{-n}(\xi)\right|$ $=\sum_{n\in\Upsilon\setminus\{0\}}\sqrt{|\mathrm{det}\;J_{\varphi_{n}}(\xi)\mathrm{det}\;J_{\varphi_{n}}(-\xi)|}e^{-\pi(\frac{5}{2})^{2}\left(\|\varphi_{n}(\xi)\|^{2}+\|\varphi_{-n}(\xi)\|^{2}\right)}$
$=K_{\xi}\max_{n\in\Upsilon\setminus\{0\}}|\mathrm{det}\;J_{\varphi_{n}}(\xi)|\sum_{n\in\Gamma_{\xi}}e^{-\pi(\frac{5}{2})^{2}\left(\|\varphi_{n}(\xi)\|^{2}+\|\varphi_{-n}(\xi)\|^{2}\right)}$
$=K_{\xi}\max_{n\in\Upsilon\setminus\{0\}}|\mathrm{det}\;J_{\varphi_{n}}(\xi)|\int_{{\mathbb{R}}^{4}}e^{-\pi(\frac{5}{2})^{2}\|u\|^{2}}{\mathrm{d}}u<\infty.$
@@
Then, Theorem 2 applies and gives a dual frame of $\{\phi_{n}\}_{n\in\Upsilon^{+}}$. The reconstruction is then given by Equation 9. ∎

### v2

@@ 4.2.2 Non-compactly supported $\boldsymbol{\widehat{\psi}}$
In this section, we assume that the support of $\widehat{\psi}$ is not compact. The following theorems state sufficient conditions for which the system
$\{T_{b_{n}k}\psi_{n}\}_{(n,k)\in\Upsilon\times{\mathbb{Z}}^{N}}\quad\textrm{ and }\quad\{T_{b_{n}k}\chi_{n}\}_{(n,k)\in\Upsilon^{+}\times{\mathbb{Z}}^{N}}$
@@
$=\sum_{k\in{\mathbb{Z}}^{N}}\left|\int_{b_{n}^{-1}\mathbb{T}}\left(\sum_{l\in{\mathbb{Z}}^{N}}\widehat{f}(\xi-b_{n}^{-1}l)\overline{\widehat{\psi}_{n}(\xi-b_{n}^{-1}l)}\right)e^{2\pi\imath\frac{1}{b_{n}^{-1}}\left(k\cdot\xi\right)}{\mathrm{d}}\xi\right|^{2},$
@@
$A=\inf_{\xi\in{\mathbb{R}}^{N}}\left(\sum_{n\in\Upsilon^{+}}\frac{1}{|b_{n}|}\left|\widehat{\chi}_{n}(\xi)\right|^{2}-\sum_{n\in\Upsilon^{+}}\sum_{k\neq 0}\frac{1}{|b_{n}|}\left|\widehat{\chi}_{n}(\xi)\widehat{\chi}_{n}(\xi-b_{n}^{-1}k)\right|\right)>0$
@@
$B=\sup_{\xi\in{\mathbb{R}}^{N}}\sum_{n\in\Upsilon^{+}}\sum_{k\in\mathbb{Z}^{N}}\frac{1}{|b_{n}|}\left|\widehat{\chi}_{n}(\xi)\widehat{\chi}_{n}(\xi-b_{n}^{-1}k)\right|<\infty,$
then the system $\{T_{b_{n}k}\chi_{n}\}_{n\in\Upsilon^{+},k\in\mathbb{Z}^{N}}$ is a frame for $\mathrm{L}^{2}({\mathbb{R}}^{N})$ with frame bounds $A$ and $B$.
@@
This results from the argumentation of the proof of Theorem 5 by replacing $\Upsilon$ by $\Upsilon^{+}$ and $\psi_{n}$ by $\chi_{n}$. ∎
@@ 5 Empirical wavelet systems from affine deformations
In this section, we consider a partition $\{\Omega_{n}\}_{n\in\Upsilon}$ for which we can find diffeomorphisms $\gamma_{n}$ that are affine functions, i.e., of the form $\gamma_{n}(\xi)=\beta_{n}(\xi-\eta_{n})$ with $\beta_{n}$ a linear function. If the construction of the empirical wavelet systems introduced in Section 3 is natural in the Fourier domain, it is also possible to build them directly in the spatial domain for affine mappings.
@@
Let $\beta:x\mapsto Ax$ be a linear function (non identically zero), with $A\in{\mathbb{R}}^{N\times N}$. The function $g$ corresponding to the inverse Fourier transform of the deformation of the Fourier transform $\widehat{f}$ of a function $f$ by $\beta$, i.e., $\widehat{g}=\widehat{f}\circ\beta$, is given by, for every $x\in{\mathbb{R}}^{N}$,
$g(x)=\frac{1}{\left|det\;A\right|}f(A^{-\intercal}x),$ (12)
@@
$g(x)=\int_{{\mathbb{R}}^{N}}\widehat{f}(\beta(\xi))e^{2\pi\imath(\xi\cdot x)}\mathrm{d}\xi$ $=\int_{\beta({\mathbb{R}}^{N})}\widehat{f}(u)e^{2\pi\imath(\beta^{-1}(u)\cdot x)}\left|\mathrm{det}\;J_{\beta^{-1}}(u)\right|\mathrm{d}u$
$=\frac{1}{\left|det\;A\right|}\int_{{\mathbb{R}}^{N}}\widehat{f}(u)e^{2\pi\imath(u\cdot\beta^{-\intercal}(x))}\mathrm{d}u$
$=\frac{1}{\left|det\;A\right|}{\mathcal{F}}^{-1}\left(\widehat{f}\right)(\beta^{-\intercal}(x))$
$=\frac{1}{\left|det\;A\right|}f(A^{-\intercal}x).$
@@
Let $\{\Omega_{n}\}_{n\in\Upsilon}$ be a partition of the Fourier domain and $\psi$ a wavelet kernel. Let assume there exists a set of linear functions $\{\beta_{n}:\xi\mapsto A_{n}\xi\}_{n\in\Upsilon}$ such that $\Lambda=\beta_{n}(\Omega_{n}-\eta_{n})$ if $\Omega_{n}$ is bounded and $\Lambda\subsetneq\beta_{n}(\Omega_{n}-\eta_{n})$ otherwise. The set of empirical wavelets, $\{\psi_{n}\}_{n\in\Upsilon}$, defined in the Fourier domain by, for every $\xi\in{\mathbb{R}}^{N}$,
$\widehat{\psi}_{n}(\xi)=\sqrt{\left|\mathrm{det}J_{\beta_{n}}(\xi-\eta_{n})\right|}\,\widehat{\psi}(\beta_{n}(\xi-\eta_{n}))=\sqrt{\left|\mathrm{det}A_{n}\right|}\,\widehat{\psi}(A_{n}(\xi-\eta_{n})),$
@@
$\psi_{n}(x)=\frac{1}{\sqrt{\left|\mathrm{det}A_{n}\right|}}\psi(A_{n}^{-\intercal}x)\,e^{2\pi\imath\,\eta_{n}x}.$
@@
$\psi_{n}(x)={\mathcal{F}}^{-1}\left(\sqrt{\left|\mathrm{det}A_{n}\right|}\widehat{\psi}\circ\beta_{n}\right)(x)\,e^{2\pi\imath\eta_{n}x}.$
By applying Lemma 1 with $\widehat{f}=\sqrt{\left|\mathrm{det}A_{n}\right|}\widehat{\psi}$ and $\beta=\beta_{n}$, we obtain the result. ∎
@@
As in Example 3.1, we consider the 1D case with $\widehat{\psi}$ of support $\Lambda$ the open interval of size $1$ centered on 00 and $\{\Omega_{n}\}_{n\in\Upsilon}$ a partition of open bounded intervals with center $\omega_{n}$. The diffeomorphism $\gamma_{n}:\xi\mapsto\frac{1}{|\Omega_{n}|}(\xi-\omega_{n})$, satisfying $\Lambda=\gamma_{n}(\Omega_{n})$, can be rewritten $\gamma_{n}(\xi)=\beta_{n}(\xi-\eta_{n})$ with $\beta_{n}:\xi\mapsto\xi/|\Omega_{n}|$ and $\eta_{n}=\omega_{n}$. The associated empirical wavelet system is therefore given in the spatial domain by, for every $x\in{\mathbb{R}}$,
$\psi_{n}(x)$ $=\sqrt{|\Omega_{n}|}\,\psi(|\Omega_{n}|x)\,e^{2\pi\imath\,\omega_{n}x}.$
@@
We can also build wavelets on a 2D dyadic tiling. We consider a separable wavelet kernel $\widehat{\psi}^{\rm 2D}(\xi_{1},\xi_{2})=\widehat{\psi}(\xi_{1})\widehat{\psi}(\xi_{2})$, where $\widehat{\psi}$ is a scaling function supported by an open interval $\Lambda$ centered in 00, and the diffeomorphism $\gamma_{j,n}:\xi\mapsto 2^{j}(\xi-\omega_{n})$, $n=1,2,3$, with $\omega_{1}=(\omega_{0},0)$, $\omega_{2}=(0,\omega_{0})$ and $\omega_{3}=(\omega_{0},\omega_{0})$. The associated translated empirical wavelet system is given in the spatial domain by, for every $(x,y)\in{\mathbb{R}}^{2}$ and $(l,m)\in{\mathbb{Z}}^{2}$,
$\psi_{j,n}((x,y)-2^{j}(l,m))$ $=\frac{1}{2^{j}}\,\psi\left(\frac{x-2^{j}l}{2^{j}}\right)\psi\left(\frac{y-2^{j}m}{2^{j}}\right).$
Finally, the transform resulting from definition 6 using this system is similar to the classic discrete wavelet transform.
The construction of a symmetric empirical wavelet system $\{\chi_{n}\}_{n\in\Upsilon}$, induced by affine diffeomorphisms, in the spatial domain stems from Proposition 2, using Equation 3.
Example 5.3.
We consider again the 1D domain. From the system $\{\psi_{n}\}_{n\in\Upsilon}$ defined as in Example 5.1, we can build a symmetric system $\{\chi_{n}\}_{n\in\Upsilon}$ as follows, for every $x\in{\mathbb{R}}$,
$\chi_{0}(x)=\sqrt{|\Omega_{0}|}\,\psi(|\Omega_{0}|x),$
@@
$\chi_{n}(x)$ $=\sqrt{\frac{|\Omega_{n}|}{2}}\psi(|\Omega_{n}|x)e^{2\pi\imath\,\omega_{n}x}+\sqrt{\frac{|\Omega_{n}|}{2}}\psi(|\Omega_{n}|x)e^{-2\pi\imath\,\omega_{n}x}$
@@ 6 Construction of empirical wavelet systems
In this section, 2D continuous empirical wavelet systems are constructed from Gabor and Shannon wavelet kernels and we examine their guarantees of reconstruction. Their practical behaviors are analyzed through numerical experiments conducted on two different (real-valued) images: a toy image of size $256\times 256$ and the classic Barbara image of size $512\times 512$. These images are shown in Figure 3 (left).
@@
To numerically construct empirical wavelet systems $\{\chi_{n}\}_{n\in\Upsilon^{+}}$ as in Equation 2, we need to compute an estimate $\breve{\gamma}_{n}$ of the diffeomorphism $\gamma_{n}$. This estimation is performed using the demons algorithm [29], which is inspired by a diffusion process and is widely used in medical image registration. This algorithm estimates a displacement field representing the desired mapping by alternating between solving the flow equations and regularization. For each region $\Omega_{n}$, the pair of smoothing parameter and number of multiresolution image pyramid levels is selected by a grid search minimizing the quadratic risk $\|\Lambda-\breve{\gamma}_{n}(\Omega_{n})\|_{2}^{2}$ on the values in $(0.3,0.35,\ldots,0.7)\times(n_{P}-2,n_{P}-1,n_{P})$, where $n_{P}$ is the highest integer such that $2^{n_{P}}$ is smaller than each dimension of the image. The number of iterations at the $n_{P}$ pyramid levels are set to $(2^{4},\ldots,2^{n_{P}+1})$ from the highest to the lowest pyramid level.
For all the numerical experiments, the symmetric empirical wavelet transform is computed as in Definition 6 with $b_{n}=1$ for every $n\in\Upsilon^{+}$, so that the results for continuous wavelet frames in Section 4 apply. To each wavelet transform coefficient ${\mathcal{E}}_{\chi}^{f}(\cdot,n)$ of an image $f$ of size $M\times M$, we associate the wavelet transform spectrum defined as $|{\mathcal{E}}_{\chi}^{f}(\cdot,n)|^{2}$ and the Fourier spectrum energy $\mathrm{E}_{n}$ of the underlying region $\Omega_{n}\cup\Omega_{-n}$ that reads
@@
The Fourier transform of the 1D Gabor wavelet kernel is given by (see [8]), for every $u\in{\mathbb{R}}$,
@@
Let assume that the diffeomorphisms $\gamma_{n}$ satisfy, for $a.e.\,\xi\in{\mathbb{R}}^{2}$, |$|\{m\in\Upsilon\mid\gamma_{m}(\xi)=\gamma_{n}(\xi)\}|\leq K_{\xi}\in\mathbb{N}$ for every $n\in\Upsilon$ and $\{|\mathrm{det}\;J_{\gamma_{n}}(\xi)|\}_{n\in\Upsilon}$ is a bounded sequence. Then, the continuous reconstruction is guaranteed for the empirical wavelet systems $\{\psi_{n}\}_{n\in\Upsilon}$ and $\{\chi_{n}\}_{n\in\Upsilon}$ induced by $\widehat{\psi}^{\mathrm{G}}\circ\gamma_{n}$, and is given by Equations 8 and 9.
@@
$\sum_{n\in\Upsilon}\left|\widehat{\psi}_{n}(\xi)\right|^{2}=\sum_{n\in\Upsilon}|\mathrm{det}\;J_{\gamma_{n}}(\xi)|\left|\widehat{\psi}(\gamma_{n}(\xi))\right|^{2}=\sum_{n\in\Upsilon}|\mathrm{det}\;J_{\gamma_{n}}(\xi)|e^{-2\pi(\frac{5}{2})^{2}\|\gamma_{n}(\xi)\|^{2}}.$
First, since $\gamma_{n}$ is a diffeomorphism, we have, for every $\xi\in{\mathbb{R}}^{2}$, $|\mathrm{det}\;J_{\gamma_{n}}(\xi)|>0$ and therefore
@@
Now, let us denote $\Gamma_{\xi}=\left\{m\in\Upsilon\mid\gamma_{m}(\xi)=\gamma_{n}(\xi)\Rightarrow n\geq m\right\}$. For $a.e.\,\xi\in{\mathbb{R}}^{2}$ and every $n\in\Gamma_{\xi}$, the $\gamma_{n}(\xi)$ are all different and the condition |$|\{m\in\Upsilon\mid\gamma_{m}(\xi)=\gamma_{n}(\xi)\}|\leq K_{\xi}\in\mathbb{N}$ means that there are at most $K_{\xi}$ integers $m\in\Upsilon$ for which the $\gamma_{m}(\xi)$ have the same value as $\gamma_{n}(\xi)$. It follows that, for $a.e.\,\xi\in{\mathbb{R}}^{2}$,
$\sum_{n\in\Upsilon}\left|\widehat{\psi}_{n}(\xi)\right|^{2}$ $\leq\max_{n\in\Upsilon}|\mathrm{det}\;J_{\gamma_{n}}(\xi)|\sum_{n\in\Upsilon}e^{-2\pi(\frac{5}{2})^{2}\|\gamma_{n}(\xi)\|^{2}}$
$\leq K_{\xi}\max_{n\in\Upsilon}|\mathrm{det}\;J_{\gamma_{n}}(\xi)|\sum_{n\in\Gamma_{\xi}}e^{-2\pi(\frac{5}{2})^{2}\|\gamma_{n}(\xi)\|^{2}}$
$\leq K_{\xi}\max_{n\in\Upsilon}|\mathrm{det}\;J_{\gamma_{n}}(\xi)|\int_{{\mathbb{R}}^{2}}e^{-2\pi(\frac{5}{2})^{2}\|u\|^{2}}\mathrm{d}u<\infty,$
where the last line comes from the fact that $\{\gamma_{n}(\xi)|n\in\Gamma_{\xi}\}$ is a sampling of ${\mathbb{R}}^{2}$ and $u\mapsto e^{-2\pi(\frac{5}{2})^{2}\|u\|^{2}}$ is a positive function. Then, Theorem 1 applies and gives a dual frame of $\{\psi_{n}\}_{n\in\Upsilon}$. This guarantees the reconstruction using Equation 8.
@@
$\sum_{n\in\Upsilon\setminus\{0\}}\left|\widehat{\psi}_{n}(\xi)\right|\left|\widehat{\psi}_{-n}(\xi)\right|$ $=\sum_{n\in\Upsilon\setminus\{0\}}\sqrt{|\mathrm{det}\;J_{\gamma_{n}}(\xi)\mathrm{det}\;J_{\gamma_{n}}(-\xi)|}e^{-\pi(\frac{5}{2})^{2}\left(\|\gamma_{n}(\xi)\|^{2}+\|\gamma_{-n}(\xi)\|^{2}\right)}$
$=K_{\xi}\max_{n\in\Upsilon\setminus\{0\}}|\mathrm{det}\;J_{\gamma_{n}}(\xi)|\sum_{n\in\Gamma_{\xi}}e^{-\pi(\frac{5}{2})^{2}\left(\|\gamma_{n}(\xi)\|^{2}+\|\gamma_{-n}(\xi)\|^{2}\right)}$
$=K_{\xi}\max_{n\in\Upsilon\setminus\{0\}}|\mathrm{det}\;J_{\gamma_{n}}(\xi)|\int_{{\mathbb{R}}^{4}}e^{-\pi(\frac{5}{2})^{2}\|u\|^{2}}{\mathrm{d}}u<\infty.$
@@
Then, Theorem 2 applies and gives a dual frame of $\{\chi_{n}\}_{n\in\Upsilon^{+}}$. The reconstruction is then given by Equation 9. ∎